Suppose a temperature of -101 degrees F was recorded in Antartica and a temperature of 127 degrees F was recorded in the Sahara Desert. How many degrees warmer is 127 degrees F than -101 degrees F?
127 degrees F is _____ degrees warmer than -101 degrees F?

Answer by richwmiller(17219)

from -101 to 0 is 101 degrees
from 0 to 127 is 127
101+127=228
127 is 228 degrees warmer than -101
another way to calculate it is
127-(-101)=127+101=228